I am choosing an integer. I then subtract 10 from the integer, take the opposite of the result, add -3 and find the opposite of the new result. My result is -3. What is the original number?
Answer by dfrazzetto(283)

Written mathematically:
-(-(x-10)-3) = -3
-(-x+10-3) = -3
-(-x+7) = -3
x-7 = -3
x = 4
CHECK:
4-10 = -6, opposite is 6
6 + (-3) = 3, opposite is -3 √